### v1

# SHaRe-SSM: An Oscillatory Spiking Neural Network for Target Variable Modeling in Long Sequences

Abstract
In recent years, with the emergence of large models, there has been a significant interest in spiking neural networks (SNNs) primarily due to their energy efficiency, multiplication-free, and sparse event-based deep learning. Similarly, state space models (SSMs) in varying designs have evolved as a powerful alternative to transformers for target modeling in long sequences, thereby overcoming the quadratic dependence on sequence length of a transformer. Inspired by this progress, we here design SHaRe-SSM (Spiking Harmonic Resonate and Fire State Space Model), for target variable modeling (including both classification and regression) for very-long-range sequences. Our second-order spiking SSM, on average, performs better than transformers or first-order SSMs while circumventing multiplication operations, making it ideal for resource-constrained applications. The proposed block consumes $73\times$ less energy than second-order ANN-based SSMs for an 18k sequence, while retaining performance. To ensure learnability over the long-range sequences, we propose exploiting the stable and efficient implementation of the dynamical system using parallel scans. Moreover, for the first time, we propose a kernel-based spiking regressor using resonate and fire neurons for very long-range sequences. Our network shows superior performance on even a 50k sequence while being significantly energy-efficient. In addition, we conducted a systematic analysis of the impact of heterogeneity, dissipation, and conservation in resonate-and-fire SSMs.
1 Introduction
Spike-based deep learning is a well-established approach for efficient Artificial Intelligence, especially due to its traits of ultra-low energy consumption and sparse computation. Whether it is digital spiking neuromorphic hardware such as Loihi loihi ; ShresthaLoihi2 , TrueNorth truenorth , or analog neuromorphic hardware such as Dynapse richter2024dynap , the spike-based computation resource utilization is significantly lower than artificial neural networks.
Presently, a significant amount of these spike-based deep learning zhou2023spikformer ; 10890026 ; lee2025spikingtransformerspatialtemporalattention ; spikingssm ; stan2024learning focus on using simple integrate and fire (IF) neurons (or their leaky version (LIF)). Although IF neurons are biologically plausible, they miss key biological characteristics such as oscillatory dynamics. Such rich dynamics are observed empirically and in biophysically detailed Hodgkin-Huxley (HH) models Hodgkin1952-sd . Although the HH model provides an accurate and detailed representation of neuronal dynamics, its high computational cost makes it impractical for large-scale AI systems. But Resonate and Fire (RF) neurons izhikevich2001resonate , though computationally equivalent to IF neurons, are expressive enough to capture such rich dynamics of biological neurons. Hence, recently, there has been a significant interest in designing RF-based spiking neural networks ShresthaLoihi2 ; higuchi2024balanced ; c-silif . To the best of our knowledge, RF neurons have not been studied in a very long sequence for target variable modeling (classification as well as regression).
For sequential data processing, transformers are the de facto standard attention . But a vanilla transformer requires quadratic computational complexity with respect to the sequence length. To obviate these limitations, methods like KV caching brandon2024reducing or test-time memory updates behrouz2024titans have been proposed. However, they do not match the computational simplicity of recurrent neural networks (RNNs). Hence, there has been significant interest in designing SSMs s4 ; s5 ; mamba , including spiking neurons stan2024learning ; spikingssm ; pspikessm , which promise to get the best of both transformers and RNNs. But the existing spiking SSMs are not optimal for target variable modeling in very long sequences.
Furthermore, as detailed in (rusch2025oscillatory, ), it’s very difficult to select a state matrix for very long interactions in a first-order SSM. To overcome these constraints, they proposed LinOSS, which is a second-order state space model rusch2025oscillatory ; dlinoss . Interestingly, LinOSS attains high classification for extremely long sequence classification and regression only with a diagonal state matrix. It achieves this by leveraging stable discretizations whose spectral radius stays close to one. LinOSS is state-of-the-art in performance on extremely long-range interactions. However, it lacks key biological features, such as spike-based communication, which has also been shown to demonstrate superior energy benefits in AI applications ShresthaLoihi2 ; Imam2020-jm .
Inspired by this progress, in this work, we propose SHaRe-SSM: Spiking Harmonic Resonate and Fire State Space Model, a second-order spike-based SSM primarily targeted for extremely long-range performance. Due to spike-based computation, SHaRe-SSM is energy efficient and hence is ideal for resource-constrained applications, such as in always-on edge AI devices. In summary, our work makes the following contributions:
  • We propose and rigorously evaluate a second-order spiking SSM: SHaRe-SSM. Noteworthily, unlike some of the previous first-order SSMs pspikessm ; spikingssm ; stan2024learning , SHaRe-SSM is completely spike-based without using non-linearities such as GeLU, GLU, or GSU.
  • As part of it, for rapid training and inference, we propose a compatible parallel scan algorithm.
  • Furthermore, we observe that SHaRE-SSM outperforms all first-order SSMs on very long sequence classification. And, SHaRE-SSM is significantly energy efficient than ANN-based second-order SSMs.
  • Also, we extend SHaRe-SSM to energy-efficient long-range regression applications and propose a convolving kernel-based regression. We observe that SHaRe-SSM outperforms first-order SSMs on a 50K sequence length regression task.
  • Finally, we study and explore the impact of discretization methods and network heterogeneities on SHaRe-SSM.
2 Related Works
State Space Models:
Structured State Space Models (SSMs), including S4, Mamba, and LRU, have demonstrated state-of-the-art performance on a wide range of sequence modeling tasks, particularly in language modeling mamba and time series forecasting rusch2025oscillatory . These models leverage the mathematical structure of state space representations to model long-range dependencies efficiently and scalably s4 ; s5 ; lru . Recent work shows that neural oscillators, as universal function approximators, naturally capture long-range temporal dependencies and often outperform traditional models cornn ; unicornn ; lanthaler2023neural . They benefit from numerically stable Implicit-Explicit (IMEX) and Implicit Euler (IM) discretizations, which better handle complex dynamics than standard Explicit Euler methods. LinOSS rusch2025oscillatory builds on this oscillator-based perspective, proposing a linear implicit oscillator-based state-space model that combines the stability of implicit solvers with the expressivity of neural oscillatory dynamics. This leads to better inductive biases for learning long sequences, especially when capturing slow or multi-scale temporal phenomena. However, LinOSS is not optimal for resource-constrained applications (such as in edge AI) due to its computational complexity.
Spiking Neural Networks:
Spiking Neural Networks (SNNs) form a biologically inspired class of neural models characterized by discrete spike-based communication. The most foundational models include the IF, LIF neurons. These have been extended to include adaptation mechanisms, such as in Adaptive LIF (ALIF), and more complex dynamics as in the RF izhikevich2001resonate and Harmonic Resonate-and-Fire (HRF) hrf neurons. Many ALIF-based recurrent models have been proposed alif ; adlif ; d-cadlif , with recent works extending adaptation to RF and HRF neurons hrf ; higuchi2024balanced , which have been evaluated on only small sequence datasets. SNNs are increasingly modeled within the SSM framework for efficient computation, especially on resource-constrained hardware. (c-silif, ) reformulated LIF and RF neurons as SSMs following (adlif, ), achieving strong results on neuromorphic benchmarks. Yet, scaling to extremely long-range sequences remains a key challenge. Recent non-adaptive, first-order SSM-based spiking models such as S4D-/S4-inspired LIF neurons stan2024learning ; pspikessm ; spikingssm and RF-based variants prf perform well on mid-range tasks (up to 16k sequence length) but fall short on longer sequences. Notably, none have been evaluated on extremely long-range time series or regression benchmarks. While an FFT-based transformer wu2025spikf has been introduced for regression, its reliance on computationally heavy FFTs limits suitability for neuromorphic deployment pspikessm . (meyer2025diagonal, ) recently validated the efficacy of an S4D-based spiking SSM by implementing it on Loihi 2. Although this design was efficient, it was evaluated on reasonably simple sCIFAR-10 datasets only.
(perez-nieves_neural_2021, ; shen2024rethinking, ) showed that in LIF/IF-based networks, neural heterogeneity—variation in neuron hyperparameters can improve classification performance. In SHaRe-SSM, we aim to explore this further by focusing on HRF neurons in a second-order SSM framework, thereby investigating the combined effects of heterogeneity and discretization. Most prior studies have relied on explicit discretization schemes, with very few exploring alternatives. A notable exception is (se-adlif, ), which demonstrated improved performance using a symplectic implicit-explicit (IMEX) scheme over conventional adaptive-LIF models adlif . We here focus on enhancing the expressivity of HRF neurons by integrating them into a fully spike-based SSM framework (SHaRe-SSM).
3 Methods
3.1 Network Description
The Resonate-and-Fire (RF) neuron, introduced by izhikevich2001resonate , better approximates the Hodgkin-Huxley (HH) model than the Leaky Integrate-and-Fire (LIF) neuron. This is because it captures complex-valued subthreshold dynamics using a 2D linear system with a complex eigenvalue, enabling membrane resonance and frequency preference. Unlike the purely real-valued LIF neuron, RF can exhibit oscillatory behavior and selective amplification of certain input frequencies, features also observed in HH neurons. Later, second-order real-valued approximations, e.g., (hrf, ; higuchi2024balanced, ), simplify RF into a real harmonic-oscillator-like form often referred to as HRF. HRF preserves oscillatory behavior via complex eigenvalues similar to the RF neuron, but without requiring initialization in complex space; The RF model’s inherent complex dynamics are a key reason it is considered more biophysically realistic. Recently, (higuchi2024balanced, ) showed RF-based models work best without resetting the membrane potential. After collating these observations, we here integrate HRF into an SSM framework and compute neuronal dynamics linearly using parallel scans, and use the spike function as an activation function to propagate spikes.
A Harmonic-Resonate-and-Fire Neuron is defined by:
$\begin{split}u^{\prime}(t)&=-\omega^{2}v(t)-2bu(t)+x(t)\\ v^{\prime}(t)&=u(t)\\ z(t)&=\Theta(u(t)-\theta)\end{split}$ (1)
where $v(t)$ denote the hidden state, $u(t)$ as the time derivative of $v(t)$, $x(t)$ is the input spikes to the neuron, $\omega\in\mathbb{R}$ is the frequency parameter and $b\in\mathbb{R}$ is the damping coefficient. Spiking Function: $\Theta\in\{0,1\}$ i.e. spikes are emitted if $u\geq\theta$.
Figure 1: Description of SHaRe-SSM model. SHaRe-SSM consists of a SHaRe-SSM neuron representing multiple states (Eq. 1) followed by a Linear layer repeated for N blocks. Spikes are encoded with a Spike Encoder and decoded using a Spike Decoder.
This HRF formulation is limited to a single neuron. We propose formulating HRF in an SSM framework (SHaRe-SSM). This approach helps in capturing better dynamics. The input spikes are multiplied by a weight matrix, and we propagate spikes from the hidden state $v(t)$. Moreover, removing the damping parameter in such second-order approximations enhances the neuron’s ability to capture long-range temporal dependencies by preserving longer oscillations rusch2025oscillatory . We define the SHaRe-SSM model by:
$\begin{split}u^{\prime}(t)&=-\Omega v(t)+Bx(t)\\ v^{\prime}(t)&=u(t)\\ z(t)&=\Theta(v(t)-\theta)\\ y(t)&=Cz(t)+Dx(t)\end{split}$ (2)
where $u(t),v(t)\in\mathbb{R}^{p}$ denote the hidden states, $y(t)\in\mathbb{R}^{h}$ the output, and $x(t)\in\mathbb{R}^{h}$ the input spike signal. The system is defined by weights $\Omega\in\mathbb{R}^{p\times p}$ which is diagonal, $B\in\mathbb{R}^{p\times h}$, $C\in\mathbb{R}^{h\times p}$, $D\in\mathbb{R}^{h}$, and an output learnable spiking parameter $\theta\in\mathbb{R}^{p}$ such that spikes are emitted if $v\geq\theta$. These thresholds ($\theta$) are learned using a step-double Gaussian surrogate gradient neftci2019surrogate ; higuchi2024balanced . Since higuchi2024balanced showed that HRF works well with no reset, we here employed the same strategy. This avoids the need for sequential processing, and the model can be implemented across time as an activation function. Figure 1 describes the SHaRe-SSM design. Each SHaRe-SSM block is comprised of SHaRe-SSM neuron, followed by a linear layer, batch normalization, and an IF neuron. We mix spikes from the previous layer and introduce dropout layers to avoid learnable thresholds from saturating early. SHaRe-SSM uses only binary spike-based communication between blocks.
3.2 Encoder and Decoder
For SHaRe-SSM, we design an encoder for encoding the input signal into spike trains (see Figure 1). Herein, the inputs every timestep are passed through a learnable linear layer. The output of this layer is passed through a batch normalization layer. Finally, an IF neuron generates spikes into the SHaRe-SSM block. Notably, this is a data-dependent trainable encoder (without requiring us to specify whether to utilize rate coding or any other encoding variants). Moreover, it utilizes neuronal heterogeneity and is instantaneous.
The decoder projects the spikes back using a linear Layer with the output dimension equal to the number of classes. For regression, we also propose convolving the output with a learnable filter.
3.3 Discretization methods
For understanding the impact of discretization on an HRF neuron (Eq. 1), we conduct a single neuron simulation (Figure 2) by injecting a unit spike input to an HRF neuron with zero phase. For the continuous case, if the damping parameter ’b’ is constrained to be non-negative, we must get stable dynamics. If set to zero, we should observe sustained oscillation. From Figure 2, we can see that for Euler Forward (Explicit), our model is diverging over time, making it unstable. IMEX discretization is overall stable and hence preserves energy. IM, although dissipative in energy, is still stable. We briefly describe IM and IMEX generalizability for long sequences in detail below.
Figure 2: Post-spike membrane potential behavior for a Harmonic RF neuron for different discretizations with no damping. The dotted red line denotes the inter-spike interval.
IMEX is promising, as (rusch2025oscillatory, ) showed that removing damping leads to a Hamiltonian system arnold1989mathematical , where energy conservation is ensured under symplectic discretizations. In addition, we examine the non-conservative but stable IM scheme, which performs well empirically. We adopt a second-order system with position-like state $u_{n}$ and velocity-like state $v_{n}$. The general form of discrete updates is denoted by:
$u_{n}$ $=u_{n-1}+\Delta t(-\Omega v_{\star}+Bx_{n}),$ (3)
$v_{n}$ $=v_{n-1}+\Delta t\,u_{n},$ (4)
$s_{n}$ $=Ms_{n-1}+F_{n},$ (5)
where the choice of $v_{\star}$ distinguishes the discretization schemes: $v_{\star}=v_{n}$ for IM (implicit), and $v_{\star}=v_{n-1}$ for IMEX (implicit-explicit).
For the IM scheme, we define rusch2025oscillatory :
$M^{\text{IM}}$ $=\begin{pmatrix}S&-S\Delta t\Omega\\ S\Delta t&S\end{pmatrix},\quad F_{n}^{\text{IM}}=\begin{pmatrix}S\Delta tBx_{n}\\ S\Delta t^{2}Bx_{n}\end{pmatrix}$,
with $S=(I+\Delta t^{2}\Omega)^{-1}$, acting as a Schur complement that suppresses high-frequency components and ensures eigenvalues of $M^{\text{IM}}$ remain bounded within the unit circle.
And, for the IMEX scheme, we use rusch2025oscillatory :
$M^{\text{IMEX}}$ $=\begin{pmatrix}I&-\Delta t\Omega\\ \Delta tI&I-\Delta t^{2}\Omega\end{pmatrix},\quad F_{n}^{\text{IMEX}}=\begin{pmatrix}\Delta tBx_{n}\\ \Delta t^{2}Bx_{n}\end{pmatrix}$.
Unlike IM, the eigenvalues of $M^{\text{IMEX}}$ lie near the unit circle, preserving oscillatory energy over long horizons. This makes IMEX especially suitable for modeling long sequences, where maintaining temporal structure is crucial.
3.4 Parallel scan in SHaRe-SSM
Parallel scans kogge1973parallel ; blelloch1990prefix exploit associativity to reduce the time complexity of recurrent computations from $\mathcal{O}(N)$ to $\mathcal{O}(\log N)$. Originally developed to accelerate RNNs, they have recently been adapted for state-space models s5 , powering modern architectures like Linear Recurrent Units (LRUs) lru and Mamba mamba . Leveraging this structure allows for richer exploration of neuronal state trajectories and improves training efficiency. We apply parallel scans to the linear computations, followed by spike function activations.
An associative binary operation similar to rusch2025oscillatory defined as:
$(a_{1},a_{2})\bullet(b_{1},b_{2})=(b_{1}\cdot a_{1},b_{1}\cdot a_{2}+b_{2}),$
where $\cdot$ represents matrix-matrix and matrix-vector multiplications. Applying a parallel scan to the input sequence $\left[(M,F_{1}),(M,F_{2}),\ldots\right]$ efficiently computes the solution to the $s_{n}=Ms_{n-1}+F_{n}$. The scan output stores $x_{n}$ in the second element of each tuple. This is made efficient by exploiting structured matrices (e.g., diagonal block $2\times 2$ matrices in $M_{IM}$ and $M_{IMEX}$), where each product can be computed in linear time with respect to the hidden dimension. We use this approach to leverage IMEX and IM discretization schemes on an HRF neuron.
Algorithm 1 describes our SHaRE-SSM implementation.
Algorithm 1 SHaRe-SSM Algorithm
0: Input sequence $x$
0:NN-blocks, spike function $\Theta$ , output sequence $o$
$x^{0}\leftarrow\textit{Encoder}(x)$ {Encode input sequence into spikes}
for $n=1,\dots,N$ do
  $z^{n}\leftarrow$ solution of HRF in (2) with input $x^{n-1}$ via parallel scan aggregated
  $z^{n}\leftarrow\Theta(y^{n}-\theta^{n}_{C})$}
  $y^{n}\leftarrow Cz^{n}+Dx^{n-1}$ {Weighted spike mixing in (2)}
  $y^{n}\leftarrow drop(\Theta(y^{n}-\theta_{D}^{n}))$
  $y^{n}\leftarrow Linear(y^{n})$
  $y^{n}\leftarrow\textit{BN}(y^{n}))$
  $y^{n}\leftarrow drop(\Theta(y^{n}-\theta^{n})))$
  $x^{n}\leftarrow cat(x^{n-1},y^{n})$
end for
$o\leftarrow\textit{Decoder}(x^{N})$ {Decode spikes}
4 Emperical Results
4.1 Long-range sequence classification
Table 1: Classification performance (mean and standard deviation across five training runs) for UEA time-series datasets using hyperparameter protocol from walker2024log . For convenience, some dataset names are abbreviated: Worms (EigenWorms), SCP1 (SelfRegulationSCP1), SCP2 (SelfRegulationSCP2), Ethanol (EthanolConcentration), Heartbeat, and Motor (MotorImagery). The top two models are bold and underlined, respectively. SNN implies a spiking neural network(Y/N).
Seq. Length/ Class 17,984/5 896/2 1,152/2 1,751/4 405/2 3,000/2
SNN Worms SCP1 SCP2 Ethanol Heartbeat Motor Avg
NRDE N $83.9\pm 7.3$ $80.9\pm 2.5$ $53.7\pm 6.9$ $25.3\pm 1.8$ $72.9\pm 4.8$ $47.0\pm 5.7$ $60.6$
NCDE N $75.0\pm 3.9$ $79.8\pm 5.6$ $53.0\pm 2.8$ $29.9\pm 6.5$ $73.9\pm 2.6$ $49.5\pm 2.8$ $60.2$
Log-NCDE N $85.6\pm 5.1$ $83.1\pm 2.8$ $53.7\pm 4.1$ $\bf 34.4\pm 6.4$ $75.2\pm 4.6$ $53.7\pm 5.3$ $64.3$
LRU N $87.8\pm 2.8$}} $82.6\pm 3.4$ $51.2\pm 3.6$ $21.5\pm 2.1$ $\bf 78.4\pm 6.7$ $48.4\pm 5.0$ $61.7$
S5 N $81.1\pm 3.7$ $\bf{89.9\pm 4.6}$} $50.5\pm 2.6$ $24.1\pm 4.3$ $77.7\pm 5.5$ $47.7\pm 5.5$ $61.8$
S6 N $85.0\pm 16.1$ $82.8\pm 2.7$ $49.9\pm 9.4$ $26.4\pm 6.4$ $76.5\pm 8.3$ $51.3\pm 4.7$ $62.0$
Mamba N $70.9\pm 15.8$ $80.7\pm 1.4$ $48.2\pm 3.9$ $27.9\pm 4.5$ $76.2\pm 3.8$ $47.7\pm 4.5$ $58.6$
LinOSS (IM) N $\bf{95.0\pm 4.4}$} $87.8\pm 2.6$}} $58.2\pm 6.9$}} $29.9\pm 0.6$}} $75.8\pm 3.7$ $60.0\pm 7.5$}} 67.8{67.8}
LinOSS (IMEX) N $80.0\pm 2.7$ $87.5\pm 4.0$}} $\bf{58.9\pm 8.1}$} $29.9\pm 1.0$ $75.5\pm 4.3$}} $57.9\pm 5.3$}} 65.0{{65.0}}
D-LinOSS N $93.9\pm 3.2$ $88.9\pm 3.0$ $58.6\pm 2.3$ $29.9\pm 0.6$ $75.8\pm 4.9$ $\bf 61.1\pm 2.0$ $\bf 68.0$
Ours (IM) Y $85.6\pm 8.1$ $82.6\pm 4.5$} $\bf 55.4\pm 6.9$} $30.1\pm 1.47$ $\bf 74.1\pm 5.9$} $\bf 59.6\pm 5.4$ $\bf 64.4$
Ours (IMEX) Y $\bf 85.6\pm 2.7$ $\bf 82.6\pm 4.2$ $53.0\pm 4.8$ $\bf 32.4\pm 4.6$ $73.2\pm 5.73$ $58.6\pm 7.3$ $64.2$
For classification experiments, we show SHaRe-SSM’s efficacy on the long-range sequential benchmark introduced in walker2024log ; rusch2025oscillatory with pre-selected random seeds for splitting the dataset (70/15/15 splits) and a fixed model hyperparameter range. It consists of the six longest sequence datasets pertaining to real-world applications from the University of East Anglia Multivariate Time Series Classification Archive (UEA-MTSCA) bagnall2018uea . To the best of our knowledge, SHaRe-SSM is the first second-order spiking SSM proposed. In addition, the recently proposed ANN-based second-order SSM (LinOSS) rusch2025oscillatory very clearly establishes its superiority over first-order SSMs. Moreover, presently proposed spiking first-order SSM’s performance pspikessm ; spikingssm ; stan2024learning is bounded by S4D’s performance. Hence, for these extremely long-range classification tasks, we compare our model performance only to the ANN-based State Space Modelss5 ; mamba ; rusch2025oscillatory (First and Second order) and Neural controlled ODEsnrde ; ncde ; walker2024log only. ANN-based LinOSS rusch2025oscillatory leverages second-order ODEs for long-range sequence learning and outperforms other SOTA models on such very long-range benchmark datasets. Similarly, SHaRe-SSM, on average, performs better than Mamba & S6 by $5.7\%$ & $2.3\%$ (See Table 1 respectively, while being only 3.6% lower than SOTA D-LinOSS. Our simplistic but fully spike-based model design (SHaRe-SSM), which leverages HRF neuronal dynamics, performs just $9.4\%$ lower than LinOSS for IM discretization on the 18k Eigenworms dataset (noteworthily (pspikessm, ) also observed a 9.4% performance drop on the Pathfinder dataset when compared to S4). Whereas for IMEX discretization, SHaRe-SSM outperforms LinOSS by $5.6\%$, while being extremely energy efficient (see Table 1). Moreover, we notice a trend that even though accuracy is the same, IMEX standard deviation is lower by $5.41\%$ and $0.32\%$ than IM for EigenWorms and SCP1, respectively. This strengthens the claim that IM introduces dissipative terms that can impact model performance. Similarly, we observe that for the Ethanol data set, the IMEX discretization on SHaRe-SSM outperforms all variants of LinOSS by $2.5\%$. While Log-NCDE outperforms our model, it comes with its own ODE computational overheads compared to LinOSS. On the 3k MotorImagery dataset, SHaRe-SSM achieves a performance of 59.6% (vs. 61.1% for D-LinOSS). Overall, our second-order spiking SSM outperforms first-order SSMs based on Mamba. Notably, a drop in performance with the gain in efficiency is ubiquitous across spiking SSMs for LRA tasks stan2024learning ; pspikessm ; spikingssm . To ensure a fair comparison, we followed the hyperparameter protocol from walker2024log ; However, constrained by computational resources and time, instead of grid search, we ran Bayesian search using Optuna akiba2019optuna for 15 runs and performed a narrow search.
4.1.1 Energy computation
Figure 3: Variation of energy computation of a LinOSS Block rusch2025oscillatory to the SHaRe-SSM Block for EigenWorms with respect to State (P), Hidden size (H).
In this section, we assess and compare the energy computation for LinOSS and SHaRe-SSM for similar hidden-size(H), state-size(P), and num-blocks(N) on the EigenWorms dataset. We estimate the theoretical energy consumption of our model based on prior works spikingssm ; pspikessm ; wu2025spikf . Accordingly, we assume that MAC and AC operations are implemented on a 45nm hardware horowitz20141 , where $E_{\text{MAC}}=4.6\text{pJ}$ and $E_{\text{AC}}=0.9\text{pJ}$. Notably, as discussed in  (spikingssm, ), the computational cost of multiplication of a floating-point weight by a binary activation number is represented as an addition-only operation. For ANNs, the theoretical energy consumption of a block $n$ is given by $4.6\text{pJ}\times\text{FLOPs}(n)$. For SNNs, the energy consumption for $n$ is given by $0.9\text{pJ}\times\text{SOPs}(n)$. Calculating theoretical energy consumption requires first calculating the synaptic operations, $\text{SOPs}(n)=f_{r}\times\text{FLOPs}(n)$, $f_{r}$ is the firing rate of the input spike train of the block/layer, $\text{FLOPs}(n)$ refers to the number of floating-point operations in layer $n$, equivalent to the number of multiply-and-accumulate (MAC) operations. SOPs denote the number of spike-based accumulate (AC) operations.
Since $\Omega$ is a diagonal matrix, it can be efficiently implemented using parallel scans with computations of order $\mathcal{O}(P\,\log(L))$, which is negligible. Also, D is $\mathcal{O}(HL)$, which is much smaller than $\mathcal{O}(LPH)$, i.e., computation for B & C matrices or even $\mathcal{O}(LH^{2})$, which is for the GLU layer. rusch2025oscillatory uses GeLU and GLU non-linearities, which we replace with a linear layer. And, yu2023metaformer described that GeLU consumes 14 FLOPs per operation. Moreover, GLU has twice as many FLOPs as a linear layer.
From Figure 3, we observe that for different state-to-hidden ratios, we get different energy numbers. For the EigenWorms dataset, the average spike rate per time step per block is $0.32$. For a lower P/H ratio, we yield high energy efficiencies ($\sim 120\times$). Even for a high ratio, our model block is $\sim 20\times$ more energy-efficient than a LinOSS Block for EigenWorms. Our best performing model (as per Table 1) is $\textbf{73.2}\times$ more energy efficient than an equivalent LinOSS model.
4.2 Regression with spikes
None of the previously proposed spiking SSMs report performance on regression tasks stan2024learning ; pspikessm ; spikingssm . In fact, most of the spiking deep learning research is focused on classification zhou2023spikformer ; higuchi2024balanced . Recently, a transformer-based method wu2025spikf has been proposed for spike-based regression of time series. But this approach is significantly limited (and deviates from present hardware-supported featuresShresthaLoihi2 ) due to its quadratic sequence length dependence bottleneck and computationally intensive FFT computations pspikessm . Unlike a transformer, SSMs have emerged as a powerful solution to the problem of such quadratic dependence on sequence length. To the best of our knowledge, for the first time, we report the regression performance of a spiking SSM for extremely long sequences (up to 50k length). Apparently, a spike-based output limits the output range for regression tasks, making it a difficult task for spiking neurons. Previously, a spike-free LI layer was shown to work for some datasets Henkes2024-zk . However, it is not scalable to the size of very long sequences due to the oversmoothing effect of the moving average. Another approach wu2025spikf translates spikes back into continuous values using a Linear layer. However, this method also leads to limited expressivity about the time steps. To mitigate these issues, we propose a kernel-based regression with a learnable kernel convolved over time, initialized with a decaying LI filter of fixed size. To find the best parameters, we apply Bayesian hyperparameter akiba2019optuna search instead of the grid search with the same range of hyperparameters for a fair comparison. Subsequently, we ran a grid search on he best kernel size to ensure our model doesn’t lose its representation power, nor oversmooths the outputs(see Table 5).
4.2.1 PPG-DaLiA Dataset Evaluation
To model very long-range interactions for regression tasks, we consider a 49920-length sequence following the processing and hyperparameter protocol as proposed in walker2024log . PPG-DalLiA is a multivariate time series regression dataset to predict heart rate using data collected from a wrist-worn device from 15 individuals for 150 minutes, sampled at 128Hz. We report the best MSE in the order of $10^{-2}$ after applying Bayesian search for 15 runs, and applied the same hyperparameters for different discretizations. From Table 2, we observe that SHaRe-SSM models consistently outperform all first-order SSM variants, even after spiking in nature. Our best-performing model is SHaRe-SSM-IMEX, which outperforms Mamba by $0.0057$ MSE, proving the energy-conserving nature of IMEX-discretization and our second-order resonating neuron representation power. Moreover, despite being energy efficient, we are only $0.0392$ MSE lower than second-order SSMs.
Table 2: Mean and Standard Deviation reported for PPG-DaLiA dataset using hyperparameter protocol from walker2024log for the mean-squared error (MSE) across five training runs. For a clear differentiation, spiking neural networks are kept separately.
Model SNN MSE $\times 10^{-2}$
NRDE N $9.90\pm 0.97$
NCDE N $13.54\pm 0.69$
Log-NCDE N $9.56\pm 0.59$}}
LRU N $12.17\pm 0.49$
S5 N $12.63\pm 1.25$
S6 N $12.88\pm 2.05$
Mamba N $10.65\pm 2.20$
LinOSS-IMEX N $7.5\pm 0.46$}}
LinOSS-IM N $6.4\pm 0.23$}}
D-LinOSS N $\bf{6.16\pm 0.73}$
Ours (IM) Y $11.27\pm 0.67$
Ours (IMEX) Y $\bf{10.08\pm 0.46}$
4.2.2 Weather Dataset Evaluation
To test our model’s performance, we follow the hyperparameter protocol of rusch2025oscillatory on a time series dataset that utilizes domain-specific preprocessing. We consider the most challenging task of 720-timestep prediction (hours) for the weather dataset from informer . We report the best Mean Absolute Error with mean and standard deviation for our model on IM and IMEX discretizations. Also, we perform Bayesian search instead of grid search with the same hyperparameter protocol and report performance for both discretizations. SHaRe-SSM outperforms ANN transformer-based modelsinformer ; logtrans ; reformer (Informer by $0.145$). RNN-based modelslstma ; lstnet such as LSTnet by $0.171$ and competitive performance with S4, with just a $0.008$ drop. From Table 3, we notice that IMEX has a deviation lower than $0.003$ compared to IM because of its energy-conserving properties.
Table 3: Mean and standard deviation (across five runs) of mean-absolute error (MAE) on the weather dataset while predicting future 720 timesteps. Hyperparameter protocol from rusch2025oscillatory
Model SNN MAE
Informer N $0.731$
Informer N $0.741$
LogTrans N $0.773$
Reformer N $1.575$
LSTMa N $1.109$
LSTnet N $0.757$
S4 N $0.578$
LinOSS-IMEX N $0.508$
LinOSS-IM N $0.528$
D-LinOSS N $0.486$
Ours (IMEX) Y $\bf{0.586\pm 0.001}$
Ours (IM) Y $0.586\pm 0.004$
4.3 Ablation Studies
4.3.1 Heterogeneity Analysis
Given the established sensitivity of SSMs to initializations, SHaRe-SSM is incorporated with heterogeneities for components - $\Omega$, B, C, D matrices and $\Delta t$, of the HRF-SSM, and the thresholds for the C ($\theta_{C}$), D ($\theta_{D}$) matrices and, the encoder ($\theta_{\text{encoder}}$). We did a detailed ablation study of the effect of each component’s heterogeneity on performance by making them homogeneous (initialized by a constant value across all neurons).
Table 4: Ablation study - Studying the effect of making each component fixed (homogeneous) on the EigenWorms dataset using IMEX with the best parameters for five seeds.(Here, $x^{*}=x^{-1/2}$)
Heterogeneity Value Accuracy
$\theta_{\text{encoder}}$ $\mathcal{U}(0,1]$ 1.0 $85.0\pm 7.2$
$\Omega$ $\mathcal{U}(0,1]$ 1.0 $82.2\pm 3.3$
B $\mathcal{U}(-{H}^{*},{H}^{*})$ 0.0 $85.0\pm 4.5$
C $\mathcal{U}(-{P}^{*},{P}^{*})$ 0.0 $85.0\pm 4.5$
D $\mathcal{N}(0,1)$ 0.0 $80.5\pm 4.6$
$\Delta t$ $\mathcal{U}(0,1]$ 1.0 $82.8\pm 2.7$
$\theta_{C}$ $\mathcal{U}(0,1]$ 1.0 $83.3\pm 3.9$
$\theta_{D}$ $\mathcal{U}(0,1]$ 1.0 $85.0\pm 4.8$
$\Omega$,B,C,D As Above As Above $85.5\pm 4.1$
$\theta_{C}$, $\theta_{D}$ As Above As Above $83.9\pm 5.1$
All As Above As Above $85.5\pm 4.1$
Only SSM Default Default $87.2\pm 3.3$
From Table 4, we observe that homogeneous initialization for neuronal thresholds leads to decreased performance. For homogeneous big step size, we see performance drop from 85.6 by 2.8% (See Table 4). Further, it drops down by 3.4% and 4.5% if the $\Omega$ matrix and D matrix are initialized with frequencies to be homogeneously kept as one and zero, respectively, showing the importance of heterogeneous states in these networks. It drops by 5% if D is initialized with zero. This implies, it is desirable to mix the HRF output with the input spikes. Homogeneity in thresholds in the Encoder layer doesn’t drop the performance much (only 0.6%), but for the thresholds of C and D, it drops by 1.7%, implying the importance of heterogeneous initialization for HRF neurons. We also noticed that complete homogenization of the model didn’t alter the mean accuracy much; rather, the deviation nearly doubled. Interestingly, if we just consider the SSM layer in a SHaRe-SSM block, we observe a boost in model performance for EigenWorms by $1.6\%$, possibly because of higher representation power without the linear layer.
4.3.2 Kernel Size for Regression
For very long-range sequences, to avoid oversmoothing, but at the same time increasing the representation power of a linear decoder, we introduce convolving a filter. This can be thought of as similar to an LI filter with learnable parameters. We investigate kernel size on the best performing model from Table 3. From Table 5, we observe that for the 50k PPG-DaLiA dataset, a kernel size of 64 outperforms the 32 and 128 filters by an error of 0.005.
Table 5: Ablation on kernel size for regression task
Model Kernel Size MSE $\times 10^{-2}$
IMEX 32 $10.62\pm 0.59$
64 $\bf{10.08\pm 0.46}$
128 $10.69\pm 0.45$
5 Discussion
In this work, we investigate the plausibility of an integration of second-order spiking neurons in a state space model (SSM) framework. As part of it, we propose an energy-efficient second-order spiking SSM: SHaRe-SSM using harmonic resonate and fire neurons. In addition, we design a learnable encoder, decoder, and a compatible parallel scan method. SHaRe-SSM is fully spike-based without any GeLU or GLU. And, SHaRe-SSM is uniquely designed to achieve energy-efficient target variable modeling on very long sequences. We evaluate SHaRe-SSM on multiple such real-world datasets, ranging from 400 to almost 18k sequence lengths for classification. On average, SHaRe-SSM outperforms all first-order SSMs while being extremely energy efficient. Furthermore, we observe that SHaRe-SSM also outperforms first-order SSMs in an energy-efficient regression task on very long sequences of length 50k. In the future, we intend to work on Intel Loihi2 ShresthaLoihi2 implementation of SHaRe-SSM for edge AI applications.
Appendix A Technical Appendix
A.1 Background Theory
To study the theoretical properties of a second-order ODE in SHaRe-SSM, we take inspiration from [21]. Our model, just like their model, can be formulated as energy-conserving and with dissipative attributes.
A.1.1 Implicit Discretization (IM):
We consider the implicit (backward Euler) discretization of a second-order system involving a position-like state $u_{n}$ and a velocity-like state $v_{n}$, similar to [21]. The implicit scheme is known to introduce additional dissipative terms, which contribute to the stability of the dynamics, particularly in the presence of stiffness.
The discretized updates are given by:
$u_{n}$ $=u_{n-1}+\Delta t\left(-\Omega v_{n}+Bx_{n}\right),$
$v_{n}$ $=v_{n-1}+\Delta t\,u_{n},$
where $\Omega$ is a diagonal matrix of oscillation frequencies and $B$ is an input projection matrix. Note that both $u_{n}$ and $v_{n}$ are evaluated at the future timestep, in contrast to explicit methods.
Letting the concatenated state be $s_{n}$, the above system can be written compactly as:
$Ms_{n}=s_{n-1}+F_{n},$
where
$M$ $=\begin{pmatrix}I&\Delta t\Omega\\ -\Delta tI&I\end{pmatrix},\qquad F_{n}=\begin{pmatrix}\Delta tBx_{n}\\ 0\end{pmatrix}$.
To obtain an explicit update rule, we algebraically solve the coupled system by introducing the matrix inverse $S=(I+\Delta t^{2}\Omega)^{-1}$. Substituting and simplifying yields:
$s_{n}=M^{\text{IM}}s_{n-1}+F_{n}^{\text{IM}},$ (6)
where
$M^{\text{IM}}$ $=\begin{pmatrix}S&-S\Delta t\Omega\\ S\Delta t&S\end{pmatrix},\qquad F_{n}^{\text{IM}}=\begin{pmatrix}S\Delta tBx_{n}\\ S\Delta t^{2}Bx_{n}\end{pmatrix}$.
This formulation highlights the stabilizing effect of the implicit method: the matrix $S=(I+\Delta t^{2}\Omega)^{-1}$ is a Schur complement that acts as a preconditioner that suppresses high-frequency components. Consequently, the eigenvalues of $M^{\text{IM}}$ remain bounded within the unit circle for a wide range of $\Delta t$, leading to improved numerical stability. The Schur complement can be computed in $\mathcal{O}(m)$ instead of the typical $\mathcal{O}(m^{3})$ operations using Gauss-Jordan elimination.
Proposition A.1.
Let $M^{\text{IM}}\in\mathbb{R}^{2p\times 2p}$ be the hidden-to-hidden weight matrix of the implicit model SHaRe-SSM-IM (6). We assume that $\Omega_{j}\geq 0$ for all diagonal elements $j=1,\dots,p$ of $\Omega$, and further that $\Delta t>0$. Then, the complex eigenvalues of $M^{\text{IM}}$ are given as,
$\lambda_{j_{1,2}}=\frac{1}{1+\Delta t^{2}\Omega_{j}}\pm\Delta t\frac{\sqrt{\Omega_{j}}}{1+\Delta t^{2}\Omega_{j}}$
with $\lambda_{j_{1}}=\overline{\lambda_{j_{2}}}$. Moreover, the spectral radius $\rho(M^{\text{IM}})$ is bounded by $1$, i.e., $|\lambda_{j}|\leq 1$ for all $j=1,\dots,p$.
Proof.
The matrix $M^{\text{IM}}\in\mathbb{R}^{2p\times 2p}$ is defined as
$M^{\text{IM}}=\begin{bmatrix}S&-\Delta t\,\Omega S\\ \Delta t\,S&S\end{bmatrix}$,
where $S=(I+\Delta t^{2}\Omega)^{-1}$, and $\Omega\in\mathbb{R}^{p\times p}$ is diagonal with non-negative entries $\Omega_{j}\geq 0$ for all $j=1,\dots,p$.
To determine the eigenvalues, we compute the characteristic polynomial:
$\det(M^{\text{IM}}-\lambda I)=\begin{vmatrix}S-\lambda I&-\Delta t\,\Omega S\\ \Delta t\,S&S-\lambda I\end{vmatrix}$.
Using block Gaussian elimination, we subtract $(\Delta t\,S)(S-\lambda I)^{-1}(-\Delta t\,\Omega S)$ from the lower-right block, giving
$\det(S-\lambda I)\cdot\det\left(S-\lambda I+\Delta t^{2}\,\Omega S^{2}(S-\lambda I)^{-1}\right).$
Since $S$ and $\Omega$ are diagonal and commute, this expression decouples elementwise. Let $s_{j}=\frac{1}{1+\Delta t^{2}\Omega_{j}}$ be the $j$-th diagonal element of $S$. Then for each $j=1,\dots,p$, the scalar characteristic equation becomes
$(s_{j}-\lambda)^{2}+\Delta t^{2}\Omega_{j}s_{j}^{2}=0.$
Solving this quadratic gives the eigenvalue pair
$\lambda_{j_{1,2}}=s_{j}\pm i\,\Delta t\,s_{j}\sqrt{\Omega_{j}},$
where $\lambda_{j_{1}}=\overline{\lambda_{j_{2}}}$, i.e., the pair are complex conjugates.
To compute their magnitude:
$|\lambda_{j_{1,2}}|^{2}=s_{j}^{2}(1+\Delta t^{2}\Omega_{j})=\frac{1}{1+\Delta t^{2}\Omega_{j}}\leq 1.$
Hence, all eigenvalues lie on or inside the unit circle in the complex plane, and the spectral radius satisfies $\rho(M^{\text{IM}})\leq 1$, as claimed. ∎
Proposition A.2.
Let $\{\lambda_{j}\}_{j=1}^{2p}$ denote the eigenvalues of the hidden-to-hidden matrix $M^{\text{IM}}\in\mathbb{R}^{2p\times 2p}$ of the SHaRe-SSM-IM model (6). Suppose the diagonal entries of $\Omega\in\mathbb{R}^{p\times p}$ are independently drawn as $\Omega_{j}\sim\mathcal{U}([0,\Omega_{\max}])$, with $\Omega_{\max}>0$. Then, the $N$-th moment of the magnitude of the eigenvalues is given by
$\mathbb{E}(|\lambda_{j}|^{N})=\frac{(1+\Delta t^{2}\Omega_{\max})^{1-\frac{N}{2}}-1}{\Delta t^{2}\Omega_{\max}(1-\frac{N}{2})},\quad\forall j=1,\dots,2p.$
Proof.
From Proposition A.1, each eigenvalue of $M^{\text{IM}}$ has magnitude
$|\lambda_{j}|=\sqrt{s_{j}}=\sqrt{\frac{1}{1+\Delta t^{2}\Omega_{j}}}.$
The NN-th moment of the magnitude is thus.
$\mathbb{E}(|\lambda_{j}|^{N})=\mathbb{E}\left[\left(\frac{1}{1+\Delta t^{2}\Omega_{j}}\right)^{\frac{N}{2}}\right].$
Applying the law of the unconscious statistician and the uniform distribution of $\Omega_{j}\sim\mathcal{U}([0,\Omega_{\max}])$, we write:
$\mathbb{E}(|\lambda_{j}|^{N})=\frac{1}{\Omega_{\max}}\int_{0}^{\Omega_{\max}}(1+\Delta t^{2}x)^{-\frac{N}{2}}\,dx.$
Substituting $u=1+\Delta t^{2}x$, so that $du=\Delta t^{2}\,dx$, the limits change from $x=0$ to $x=\Omega_{\max}$, corresponding to $u=1$ to $u=1+\Delta t^{2}\Omega_{\max}$. The integral becomes:
$\mathbb{E}(|\lambda_{j}|^{N})=\frac{1}{\Delta t^{2}\Omega_{\max}}\int_{1}^{1+\Delta t^{2}\Omega_{\max}}u^{-\frac{N}{2}}\,du.$
This evaluates to
$\mathbb{E}(|\lambda_{j}|^{N})=\frac{(1+\Delta t^{2}\Omega_{\max})^{1-\frac{N}{2}}-1}{\Delta t^{2}\Omega_{\max}(1-\frac{N}{2})},$
We can observe from proposition A.2 that even though the spectral radius of eigenvalues is smaller than one (proposition A.1), it is large enough to capture long-range dependencies even for very long-range sequences, even for $\Omega_{\max}=1,\Delta t=1$. Hence, we initialize $\Omega_{j}\sim\mathcal{U}([0,1]),\Delta t_{j}\sim\mathcal{U}([0,1])$
A.1.2 Implicit-Explicit Discretization (IMEX):
We also utilize an implicit-explicit (IMEX) scheme for discretizing the second-order harmonic oscillator system, similar to [21]. IMEX methods treat the stiff terms implicitly and the non-stiff or input terms explicitly, resulting in a balanced scheme that enables stable yet undamped oscillations. As shown in [21], such schemes preserve the total energy of the system and therefore are particularly well-suited for learning long-range sequential patterns without introducing artificial dissipation.
The update equations under the IMEX discretization are given by:
$u_{n}$ $=u_{n-1}+\Delta t\left(-\Omega v_{n-1}+Bx_{n}\right),$
$v_{n}$ $=v_{n-1}+\Delta t\,u_{n},$
where the velocity update depends implicitly on the newly computed $u_{n}$, while the force term $\Omega v_{n-1}+Bx_{n}$ is evaluated using previous state values.
Defining the state vector as $s_{n}$, we can rewrite the update in matrix form:
$Ms_{n}=M_{1}s_{n-1}+F_{n},$
where the matrices $M$, $M_{1}$, and input vector $F_{n}$ are:
$M$ $=\begin{pmatrix}I&0\\ -\Delta tI&I\end{pmatrix},M_{1}=\begin{pmatrix}I&-\Delta t\Omega\\ 0&I\end{pmatrix},F_{n}=\begin{pmatrix}\Delta tBx_{n}\\ 0\end{pmatrix}$.
Multiplying both sides by $M^{-1}$ yields the closed-form update:
$s_{n}=M^{\text{IMEX}}s_{n-1}+F_{n}^{\text{IMEX}},$ (7)
where the transition matrix and input vector are given by:
$M^{\text{IMEX}}$ $=\begin{pmatrix}I&-\Delta t\Omega\\ \Delta tI&I-\Delta t^{2}\Omega\end{pmatrix},\qquad F_{n}^{\text{IMEX}}=\begin{pmatrix}\Delta tBx_{n}\\ \Delta t^{2}Bx_{n}\end{pmatrix}$.
Proposition A.3.
Let $M^{\text{IMEX}}\in\mathbb{R}^{2p\times 2p}$ be the hidden-to-hidden weight matrix of the implicit-explicit model SHaRe-SSM-IMEX (7). Suppose that $\Omega\in\mathbb{R}^{p\times p}$ is diagonal with strictly positive entries $\Omega_{j}>0$ for all $j=1,\dots,p$, and that the time step satisfies $0<\Delta t\leq\max_{j}\left(\frac{2}{\sqrt{\Omega_{j}}}\right)$. Then, the eigenvalues of $M^{\text{IMEX}}$ are given by
$\lambda_{j_{1,2}}=\frac{1}{2}(2-\Delta t^{2}\Omega_{j})\pm\frac{1}{2}\sqrt{\Delta t^{2}\Omega_{j}(4-\Delta t^{2}\Omega_{j})},$
with $\lambda_{j_{1}}=\overline{\lambda_{j_{2}}}$. Moreover, the eigenvalues lie on the complex unit circle, i.e., $|\lambda_{j}|=1,\quad\forall j=1,\dots,p$.
Proof.
The matrix $M^{\text{IMEX}}\in\mathbb{R}^{2p\times 2p}$ has the block form
$M^{\text{IMEX}}=\begin{bmatrix}I&-\Delta t\,\Omega\\ \Delta t\,I&I\end{bmatrix}\begin{bmatrix}(I+\Delta t^{2}\Omega)^{-1}&0\\ 0&(I+\Delta t^{2}\Omega)^{-1}\end{bmatrix}$,
So the effective system matrix becomes
$M^{\text{IMEX}}=\begin{bmatrix}S&-\Delta t\,\Omega S\\ \Delta t\,S&S\end{bmatrix},\quad\text{where }S=(I+\Delta t^{2}\Omega)^{-1}.$
We analyze the characteristic polynomial:
$\det(M^{\text{IMEX}}-\lambda I)=\begin{vmatrix}S-\lambda I&-\Delta t\,\Omega S\\ \Delta t\,S&S-\lambda I\end{vmatrix}$.
Using block elimination, we simplify:
$\det(S-\lambda I)^{2}+\Delta t^{2}\Omega S^{2}.$
Since all matrices are diagonal, the problem decouples elementwise. Let $s_{j}=\frac{1}{1+\Delta t^{2}\Omega_{j}}$ for each $j=1,\dots,p$. Then, for each $j$, the characteristic polynomial becomes
$\lambda^{2}-(2s_{j})\lambda+(s_{j}^{2}+\Delta t^{2}\Omega_{j}s_{j}^{2})=\lambda^{2}-(2-\Delta t^{2}\Omega_{j})\lambda+1=0.$
Solving this gives the eigenvalue pair
$\lambda_{j_{1,2}}=\frac{1}{2}(2-\Delta t^{2}\Omega_{j})\pm\frac{1}{2}\sqrt{\Delta t^{2}\Omega_{j}(4-\Delta t^{2}\Omega_{j})}.$
To show $|\lambda_{j_{1,2}}|=1$, we consider two cases:
  1. 1.
    If $\Delta t^{2}\Omega_{j}=4$, then the square root vanishes and
    $\lambda_{j_{1}}=\lambda_{j_{2}}=-1,\quad|\lambda_{j_{1,2}}|=1.$
  2. 2.
    If $\Delta t^{2}\Omega_{j}<4$, then the eigenvalues are complex conjugates. Their squared magnitude is
    $|\lambda_{j_{1,2}}|^{2}$ $=\left(\frac{2-\Delta t^{2}\Omega_{j}}{2}\right)^{2}+\left(\frac{1}{2}\sqrt{\Delta t^{2}\Omega_{j}(4-\Delta t^{2}\Omega_{j})}\right)^{2}$
    $=\frac{(2-\Delta t^{2}\Omega_{j})^{2}+\Delta t^{2}\Omega_{j}(4-\Delta t^{2}\Omega_{j})}{4}$
    $=\frac{4-4\Delta t^{2}\Omega_{j}+\Delta t^{4}\Omega_{j}^{2}+4\Delta t^{2}\Omega_{j}-\Delta t^{4}\Omega_{j}^{2}}{4}=1.$
Thus, in both cases, $|\lambda_{j_{1,2}}|=1$, completing the proof. ∎